OpenStudy (anonymous):

Find the equation of the line with slope -3 containing the point (-1, 5) in both point-slope form and slope-intercept form

OpenStudy (anonymous):

Well you know point-slope to be: \[y-y _{1}=m(x-x _{1})\]And Slope intercept form: \[y=mx+b\]So for point slope you can plug in m as -3 and (-1,5) as \[(x _{1},y _{1})\] Then solve, for slope intercept form, you plug in (-1, 5) and m=-3 and solve for b and then rewrite it as y=mx+b. Hope it helps :)

OpenStudy (anonymous):

when solving for the slope intercept i know the formula is y=mx+b, but is b 5 or would it be 2 from when you solve for the point slope...the point slope i got is y=-3x+2 so does b =2. in this question would the point slope be y=-3x+5 or y=-3x+2?

OpenStudy (anonymous):

Well the coordinate y (5) is not b, remember b is the y-intercept. So, y=mx+b 5=(-3)(-1)+b solve for b here 5=3 + b 2= b Thus, y=-3x+2 just like the point-slope.

OpenStudy (anonymous):

ok, great thank you so much

OpenStudy (anonymous):

Yep, Good Luck :)